a.Suppose the total cost for various levels of output for a competitive firm are given in the table below:

Q

TC

0

10

1

12

2

15

3

19

4

24

5

30

6

37

7

46

8

55

9

65

If the market price is $8, how many units should the firm produce to maximize profit?

6

8

5

7

b,

A firm in a competitive market has the following cost structure:

Output

Total Cost

0

$5

1

$10

2

$12

3

$15

4

$24

5

$40

If the market price is $4, this firm will

shut down in the short run and exit in the long run.

produce three units in the short run and exit in the long run.

produce four units in the short run and exit in the long run.

produce two units in the short run and exit in the long run.
Answer #1

_(% Теме 10 19 24 10 (2-0) - 2 (15-12) ₂3 (19-15) = 4 (24-19) = 5 (30-24)= 6 (37-30) = 7 (46-37) = 9 (55-46) -9 (85-55)=10. 36 output Total Cast Total fixed cost Total vast. ble lest AVC MC 5 116-5)=5 lo 12 10 15 3.5 (12-10)=2 3:33K15-12),3 (24-15)=9